OpenStudy (anonymous):

Okay here is the deal, I am studying to test out of two math classes, I need someone to show step by step how to Sum of products in trig and Cos laws I have problems this not a the test. i am only studying. I will be attending IU east on-line this summer studying math. Please show me how to do it not give me answers only thanks.

OpenStudy (anonymous):

here is the problem cos6x+cos2x=0

OpenStudy (anonymous):

sin2xcos2x-cos2x=0

OpenStudy (inkyvoyd):

two tips that are meant as general self study guidelines: sign up for a free account on wolfram alpha, and use step-by-step solution after entering equations. Use resources such as Khan Academy, Purple Math (I believe), Wikipedia, and SOS math (I believe). I think paul's online math notes has good info too. Note that I have given you 7ish names and no links - trust me, these worked for me, especially the first, if you do not have openstudy to help you. Good luck, and have fun!

OpenStudy (anonymous):

Youtube also has some good math videos. Just type in sum and difference trig identity and you should find what you are you looking for.

OpenStudy (anonymous):

I know you plug in for this I am just not how?

OpenStudy (anonymous):

so we use this form in cosucosv=1/2{cosu-v)+cos(u+v0} so now we plug in?

OpenStudy (anonymous):

Cos u+Cos v =2Cos((u+v)/2)Cos((u-v)/2)

OpenStudy (anonymous):

So now just make the appropriate substitutions for u and v.

OpenStudy (anonymous):

when plugging in for U and v does the 1/2 mean a power?

OpenStudy (anonymous):

No, it's multiplying the sum and difference of the two angles by 1/2.

OpenStudy (anonymous):

cosucosv=1/2 cos 6x+cos2x=0

OpenStudy (anonymous):

is this right so far?

OpenStudy (anonymous):

Actually, it should be 2cos4xcos2x=0

OpenStudy (anonymous):

okay that is what I need help with?

OpenStudy (anonymous):

1/2 of 6 is only 3??

OpenStudy (anonymous):

Sum of 6x and 2x is 8x divided by 2 is 4x. Difference of 6x and 2x is 4x divided by 2 is 2x.

OpenStudy (anonymous):

You do not take halfe of each angle you take half of their sum and then half of their difference.

OpenStudy (anonymous):

so you add 6+2= 8 and than divide the sums by 2?

OpenStudy (anonymous):

Yes, and then you do it again only use their difference.